EVA Report – April 14th

EVA Report

EVA #9 – 14Apr2018 – Sol 6

Crew members: Richard Blakeman, Ashok Narayanamoorthi

Objectives:

To explore the Kissing Camel Ridge – West region. Conducted
reconnaissance and survey mission of the geology and communications capability with primary habitat. Continued the mission to explore Zubrin’s Head.

Operations:
0800-1100

Narrative:
Two person EVA to explore a new site and further explore one already seen. Going over sites we’ve already been to always reveals something new and exciting.

Vehicles
Blue ATVs #1 and #3
_______________

EVA Report

EVA #10 – 14Apr2018 – Sol 6

Crew members: Joe Dituri, Andreea Radulescu, Victoria Varone
Objectives:
– Explore Road 1104 and the surrounding area
– Hab skirt removal upon return

Operations:
1500 – 1800

Narrative:
Three person team explored the road and stopped several times to explore the interesting structures. Some of the hiking was difficult which assisted in the fogging study.

Vehicles
Red ATV, Blue ATV #2, Curiosity Rover

Journalist Report – April 14th

Tomorrow we burn the dead horse…

In Navy tradition when a sailor was hired for a cruise the captain would pay him half of his wages upfront and tell him to go mind his p’s and q’s…. minding his p’s and q’s meant that he should pay his bar tab for his pints and his quarts.

As today is the start of our approximate midway point…tomorrow we are working for ” pay ” as opposed to worling for free…;*) We celebrate with festivities. In old days they would burn a wooden horse as a gesture of half way complete.

Today we had a great breakfast followed by exploration EVAs which further the science we continue to work. (Studies about weight loss, fogging and muscle loss)

The afternoon EVA had a 3 person party come across a cow in the road. The commander…in the face of a forced vegetarian life style… was caught salivating over the meat…but alas the cow moved on along with his hopes of a T-bone steak.

We had a terrific meal of spaghetti and each shared a tequila story…you all know you have one…;*).

The crew never went through the forming or storming stages since we are well versed and acquainted….we jumped right into norming and quicky progressed into PERFORMING. This well oiled machine is going to take a half day safety stand down and eat Chocolate and banana pancakes…then do a little medical training and then wrap up with a short ATV tow exercise. Shouls be back in the house at 3 pm and do a MOVIE night….Doc is going to try to pop dehydrated corn for popcorn….;*)

Warm Regards,

Joseph Dituri, PhD(c)
CDR, U.S. Navy Diving Officer (ret)

Director – IBUM

Sol Summary – April 14th

Crew 192 Sol Summary Report 14Apr2018

Sol 6

Summary Title: Here Comes the Sun

Author: Victoria Varone

Mission Status: The crew is now halfway through their mission! We were very excited to get back to EVAs after the last 2-3 days of strong winds and cold temperatures.

Sol Activity Summary: Today consisted of two really great EVAs for Crew 192. The weather started off cool and sunny and gradually became warmer, remaining clear and beautiful all day.

At the end of the second EVA, we started work on a hab repair request to remove the skirting from all the way around the hab. Some of it has been removed and the rest will be worked on during our second week in sim.

We ended up the day with an Italian meal of pasta and salad. Ciao bella!

Look Ahead Plan: Some stargazing at night (we’ll be using some red astronomy lights for night vision), a half-day’s break tomorrow plus a medical simulation EVA, some more planning for hab repairs and EVAs for next week.

Anomalies in Work: None

Weather: Warm and sunny.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Report, EVA Request Journalist Report, Photo Report

Support Requested: See Operations Report

Maintenance request: See Operations Report

Operations Report – April 14th

Crew 192 Operations Report 12Apr2018

SOL 6

Name of person filing report: Victoria Varone

Non-nominal systems: Internet. WiFi was down for most of the day, every member of the crew was getting a “WiFi Connected, but No Internet” message. It’s holding now after being reset but we will keep monitoring it.

Generator (hours run): 12 hours

Solar ‒ SOC 84%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas. We have no knowledge of the gas or that 5 gallon container.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 850 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133477.4

Toilet tank emptied: No. Last emptied April 11, 2018.

ATV’s Used: Red ATV, Blue ATVs #1, #2, and #2

Oil Added? No but checked and OK.

ATV Fuel Used: 20%

# Hours the ATVs were Used today: 6

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 25.9
Beginning charge: 100%
Ending charge: 91%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well and having a great time going on EVAs and cooking together.

Summary of internet: 697MB at 2040.

Summary of suits and radios: Suits all charging, 3 one-piece suits and 2 two-piece suits used. Radios working fine.

Summary of Hab operations: Two successful EVAs, inspected and tested Musk Observatory (first day of acceptable weather), lots of cooking.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank remains in the RAM for additional cleaning.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Spacesuit 9’s ventilation fan is not working. It worked during our first EVA but is not charging.

Greenhab Report – April 14th

GreenHab Report
Name: Andreea Radulescu
Date: 14 Apr 2018

Environmental control:

Ambient with door open
Shade cloth: on
Fan: not on.

Average temperatures: 21.7℃
Low 13.7℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: Picked some tomatoes, pepper and red lettuce.

Daily water usage for crops: 8 gal

Time(s) of watering for crops: Twice: In the morning around 9:15am and afternoon 6:30pm

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: Picked some tomatoes, green pepper and red lettuce for the team. Also I have added MiracleGro when watering the plants.

Support/supplies needed: None

Commander Report

Title…SCRAMBLE DOG IS A STEELY EYED MISSLE MAN.

Great day of training / exploring.
Very productive EVAs continuing both our science (weight, muscle loss, suit fogging…) along with the work for Mars society…cleaning the water tank and removing the external skirt on the HAB.

The HAB skirt haa been removed “areas outside” our habitat. We will send a team and remove the skirt from the “inside areas” …for the purposes of our SIM ….at a later time and place on trailer along with other pieces. We cannot fit in area under fromt stairs in EVA suits so that area remains skirted.

Water tank needed some time to air out so doctor removed the cap and inspected. Intend to clean and put back on trailer early next week ready for service.

IMPRESSED: I continue to be impressed with the XO. He continues to be the voice of reason and chief moral officer. BZ Richard Blakeman.

LEARNED: Being a vegetarian by virtue of others around you being vegetarians…IS HARD. Yogi needs meat. Saw a cow on EVA….mmmmmm

POA&M….if approved by CAPCOM …would love to execute plan SCRAMBLE DOG.
Scramble dog consists of a half day “sleep in” partial day of rest for the crew ….XO and I are gonna make banana chocolate chip pancakes for the crew for beeakfast!!!!
Then… after noon …
medical training (short OJT by doctor) and a towing ATV exercise followed by movie night and an attempt at making dehydrated corn popcorn. Why NOT….;*)

Warm Regards,

Joseph Dituri, PhD(c)
CDR, U.S. Navy Diving Officer (ret)

Director – IBUM